Accelerator Update for March 3, 2008 - March 5, 2008
Monday March 3
The day shift began with the Tevatron (TeV) off and being prepared for access, with no Recycler stash, with the Antiproton Source in an AP0 access for lithium lens work, with the Main Injector (MI), Linac, and Booster operating normally, with the Collider Detector at Fermilab (CDF) on access, with DZero (D0) waiting for the next store, with NuMI off for horn repair, and with the SciBooNE / MiniBooNE (BNB) experiments taking beam.
At 9:39 AM, power went out in part of the TeV sector A0 tunnel. Electricians responded and restored the power.
At 9:50 AM, a Booster RF station (BRF9) tripped off. It reset.
Operations prepared the Pbar rings for access at 10:22 AM.
A Linac expert began a LRF5 study at 10:24 AM.
The vacuum technicians came out of F sector at 10:28 AM.
The LRF5 study ended at 11:26 AM.
At 11:29 AM, Operations reported that they could not turn on MIRF5 & 14. Experts investigated and had the stations on 20 minutes later.
The Neutron Therapy Facility (NTF) reported at 11:47 AM that they had completed treatment on their last patient for the day.
Operations established beam to BNB at 12:29 PM.
At 12:39 PM, the Linac 400 MeV area tripped off for the first of seven times in a little over two hours. It reset.
Pbar experts began to pulse the lithium lens at 2:26 PM.
Operations put the TeV through a dry squeeze at 2:46 PM.
FESS electricians reported that they had replaced two CDF breakers by 3:45 PM.
Recycler experts reported at 3:53 PM that they were conducting studies.
Pbar experts reported at 4:29 PM that they were rebuilding the vault at AP0.
Operations reported at 5:34 PM that they were sending beam to Pbar.
Operations put the TeV through a wet squeeze at 5:49 PM.
At 6:09 PM, the lithium lens tripped off. Pbar experts investigated and reported that the new lens had a water system breach. They will begin repairs on Tuesday morning.
At 6:42 PM, Operations reported that they were having trouble keeping beam in MI for shots of beam to the TeV. Experts were investigating.
At 7:39 PM, Operations reported that while setting up for a continuous injection of protons into the TeV, they had injected beam on top of existing beam and caused a TeV quench at sectors F2 and F3.
At 8:07 PM, BRF10 tripped off. It reset.
Operations put the TeV through a dry squeeze at 9:33 PM.
Operations reported at 11:48 PM that they were working with experts to resolve a MI MECAR problem.
Tuesday March 4
The midnight shift began with TeV and Recycler experts conducting studies, with MI having MECAR problems, with no stacking due to a lithium lens problem, with NuMI off for horn repair, and with BNB taking beam.
Operations reported at 12:31 AM that with the help of an expert, operators rebooted MECAR and then changed the supercycle length. This allowed MI to resume sending protons to the TeV.
Operations put the TeV through a dry squeeze at 1:11 AM.
At 1:14 AM, a TeV kicker (AAK1) prefired and caused beam to abort.
At 1:19 AM, the BNB beam permit tripped off when a number of power supplies tripped off. Everything reset.
At 1:57 AM, AAK1 prefired again and aborted beam. Operations turned off the kickers.
Operations put the TeV through a dry squeeze at 1:58 AM.
A Recycler expert completed a study at 2:20 AM.
Operations put the TeV through a wet squeeze at 2:30 AM.
At 3:04 AM, LRF5 tripped off. It reset.
At 3:18 AM, LRF5 tripped off. It reset.
At 6 AM, Operations turned off the MI beam and prepared MI for access to replace a quadrupole magnet (Q603) that had an LCW leak.
At 6:46 AM, BRF18 tripped off. It reset.
Operations prepared CDF and D0 for an access at 6:50 AM.
Pbar experts resumed work on the AP0 lithium lens around 7 AM.
Operations prepared the Pbar rings for access at 8:39 AM.
At 8:50 AM, the Booster West RF Watchdog tripped on a Power Amplifier LCW leak. Experts investigated.
At 9:49 AM, a Linac Klystron (KRF1) tripped off. It reset.
Kicker experts began work on AAK1 at 10:21 AM.
At 10:31 AM, the MI-60 interlocks dropped. Interlock experts found and repaired a failed door switch.
Operations reported at 11:40 AM that mechanical techs had removed the bad MI Q603 magnet.
Operations began preparing the TeV for access at 10:37 AM.
Operators search and secured MI-10 and MI-60 at 11:49 AM.
Linac experts reported at 12:02 PM that they had finished their LRF5 study.
Booster experts began a study at 1:03 PM.
At 1:26 PM, MI experts found another LCW leak on a quadrupole in the 208 area. They began repairs.
TeV experts reported at 2:02 PM that the B3 dump switch had been replaced.
Operations reported at 2:04 PM that TeV experts had put the antiproton kickers in conditioning mode.
At 2:18 PM, LRF5 tripped off. It reset.
TeV experts reported at 3:04 PM that the circuit supplying power to the two D0 skew quadrupole magnets had been split so that each magnet was powered separately.
Interlock techs accessed Linac and performed safety system tests at 3:46 PM.
D0 came out of access at 4:04 PM.
Operations search and secured CDF at 4:53 PM.
MI reported at 4:59 PM that the LCW leak repair on Q208 was finished.
Operations established beam to the Booster dump at 6:16 PM.
A Pbar expert began pulsing the lithium lens at 6:52 PM.
At 7:02 PM, KRF3 tripped off. It reset.
At 7:05, a Booster vacuum pump (IPRF23) tripped of and caused BRF13 and 14 to trip. Everything reset.
At 7:13 PM, A Booster vacuum pump (IPRF17) tripped off and caused BRF7 and 8 to trip. Everything reset.
The MI access ended at 7:27 PM.
Operations put the TeV through a dry squeeze at 7:43 PM.
Operations established beam to BNB at 8:25 PM.
At 9:04 PM, the TeV beam aborted. Experts investigated and found a 480-volt disconnect fuse that had failed. They replaced the failed fuse and the other two in the set.
At 9:06 PM, BRF11 tripped off. It reset.
At 9:07 PM, LRF5 tripped off and would reset, but then trip off. Operators contacted experts and then got the station back up about 40 minutes later.
Operations reported at 9:13 PM that they were waiting for the MI vacuum to stabilize before they would send beam to MI.
At 9:27 PM, the 400 MeV area tripped off for the first of two times in seven minutes due to bad beam positions. It reset.
A Recycler expert began a study at 10:39 PM.
Operations established beam to MI at 10:43 PM.
At 10:44 PM, LRF5 tripped off. It reset.
A Pbar expert had Q916 back online by 11:09 PM and Pbar began stacking.
Operations injected protons into the TeV at 11:15 PM.
Operations put the TeV through a wet squeeze at 11:33 PM.
At 11:37 PM, the lithium lens tripped off on a “Lens LCW Return Conductivity High” indication. This was the same trip indication that the last lens replacement had shown when it tripped. An expert investigated.
Pbar resumed stacking at 11:39 PM.
At 11:41 PM, the lithium lens tripped off. It reset.
At 11:43 PM, the lithium lens tripped off. It reset.
At 11:52 PM, the lithium lens tripped off. It reset.
At 11:56 PM, the lithium lens tripped off. Experts investigated.
Operators aborted the TeV beam at 11:48 PM. It wasn’t a very clean abort. Operations contacted an expert.
Wednesday March 5
The midnight shift began a TeV expert conducting studies, with Pbar off due to lens problems, with NuMI off for horn repair, and with BNB taking beam.
Operations put Pbar into standby at 12:43 AM. Lithium lens work will resume later in the morning.
The Recycler study ended at 12:51 AM.
At 1:49 AM, KRF3 tripped off. It reset.
At 2:46 AM, LRF5 tripped off. It reset.
The TEV studies ended at 4:50 AM.
At 6:26 AM, KRF3 tripped off. It reset.
At 6:29 AM, two Booster vacuum valves closed and caused a power supply (GMPS) to trip. Everything reset.
The Plans for Wednesday
The plans for today are running tests on the lithium lens. MI experts will conduct studies.
Comments and Suggestions
|last modified 03/05/2008 email Fermilab|
|Security, Privacy, Legal|